# Dynamic models of gentrification

### v1

@@ 5 Discussion
We find that even a small proportion of high-income agents (5% of the population) significantly affect the dynamics of gentrification. When high-income agents do not move, even if still present in the model city, no gentrification is detected, and the model does not converge within the imposed limit of 300 steps. However, introducing even a low probability for the movement of high-income agents ($p_{H}$ = 1%) leads to model convergence in approximately 150 steps, while 40% of the city neighbourhoods experience at least one gentrification wave.
@@
Our measures, $\mathcal{G}_{count}$ and $\mathcal{G}_{net}$, represent two distinct approaches to quantify gentrification. The count-based measure $\mathcal{G}_{count}$ evaluates the concentration of agents of the three types in each neighborhood over a time window $\Delta$. A significance threshold ($n_{H,M}^{*}$) is needed to detect when middle- and high-income agents are over-represented and, therefore, define the binarised version of the count-based measure, $\mathcal{G}_{bin}$. Such measure thus captures neighborhood transitions from under- to over-representation of middle- and high-income agents, indicating the completion of gentrification. The network-based measure $\mathcal{G}_{net}$ tracks the net inflow of middle- and high-income agents and the simultaneous outflow of lower-income agents over $\Delta$. This measure is thus rooted in temporal network analysis, where the existence of structures [47, 48, 49] in the networks under study is related to the simultaneity of the interactions (edges) between pairs or groups of nodes. Furthermore, this network-based approach avoids more or less arbitrary thresholds and aligns with available commuting flow data [35, 36, 37, 38], which could serve as a proxy given the absence of residential relocation records. $\mathcal{G}_{net}$ helps identifying early signs of gentrification by tracking peaks or plateaus in its trajectory, correlating higher-income resident influx with lower-income displacement. The cross-correlation computed between the count-based an the network-based measures highlights how peaks in $\mathcal{G}_{net}$ are consistently and significantly observed in advance with respect to $\mathcal{G}_{count}$. The analysis of the randomised version of our model, in Supplementary Figure S2, shows that while gentrification events caused by random agent-relocations are still detected, no significant cross-correlation exists between our two measures, emphasising the importance of the agents’ decision-making rules in our model for predicting transitions. These results demonstrate that our minimal model captures the essential features needed to reproduce gentrification: a heavy-tailed income distribution, few income classes, and distinct relocation strategies – notably the profit-driven behavior of high-income agents. Moreover, our network-based measure $\mathcal{G}_{net}$ enables earlier detection of gentrification compared to count-based metrics, potentially aiding policymakers in preventing low-income displacement.
@@
While our model provides valuable insights into gentrification dynamics, it has limitations that future research could address. The constant population size and static income-group assignments could be expanded to incorporate population growth and inter-city migration flows, potentially using a network-based approach to disentangle endogenous and exogenous causes of gentrification [50]. Multiple property ownership per agent could be introduced to model wealth concentration and short-term rental effects. The grid-based urban representation could be enhanced with more complex geographical features, although our results hold for both 7$\times$7 and 9$\times$9 grids (Supplementary Information 5-7). Moreover, simulating policy interventions [51] could provide insights for urban planners, particularly regarding density restrictions given our findings on city density and gentrification probability. Finally, future research In conclusion, our agent-based model and novel quantitative measures offer a powerful framework for understanding and predicting gentrification processes. This approach not only advances our theoretical understanding of gentrification but also provides quantitative what-if tools for early detection and potential mitigation of its effects in real-world urban environments.
Methods
Agents’ income.
@@ Methods
Gentrification: peaks, shifts and aggregate measure
@@
$T_{shift}^{i}=\{t:\mathcal{G}_{bin}^{i}(t,\Delta)=1\wedge\mathcal{G}_{bin}^{i}(t-1,\Delta)=0\}$ (14)
@@
$T_{peak}^{i}=\left\{t:\left[\begin{array}[]{c}(\mathcal{G}_{net}^{i}(t,\Delta)>\mathcal{G}_{net}^{i}(t-1,\Delta)\wedge\mathcal{G}_{net}^{i}(t,\Delta)>\mathcal{G}_{net}^{i}(t+1,\Delta))\\ \vee\\ (\mathcal{G}_{net}^{i}(t,\Delta)>\mathcal{G}_{net}^{i}(t-1,\Delta)\wedge\mathcal{G}_{net}^{i}(t,\Delta)=\mathcal{G}_{net}^{i}(t+1,\Delta))\end{array}\right]\right\}$ (15)
@@
$\mathcal{G}_{bin}^{city}=\frac{\sum_{i=1}^{N_{\text{cells}}}\mathbf{1}(|T_{shift}^{i}|>0)}{N_{\text{cells}}}\times 100\quad,\quad\mathcal{G}_{net}^{city}=\frac{\sum_{i=1}^{N_{\text{cells}}}\mathbf{1}(|T_{peak}^{i}|>0)}{N_{\text{cells}}}\times 100$ (16)
@@
Lagged cross-correlation
We compute the mean cross-correlation $\langle R\rangle$ by calculating the lagged cross-correlation between pairs of time series of $\mathcal{G}^{i}_{net}(t,\Delta)$ and $\mathcal{G}^{i}_{bin}(t,\Delta)$ corresponding to each cell on the grid-view of a model city, for several values of the lag $\tau$, and then averaging over all cells. To compute the cross-correlation $R^{i}$ between the two time series for a cell $i$, we transform the two time series into two binary vectors $\widetilde{\mathcal{G}}_{net}^{i}$ and $\widetilde{\mathcal{G}}_{bin}^{i}$ of length $T$, where the $t_{k}$-th entry is $1$ if the original corresponding time series has a peak or 0-1 transition at time $t_{k}\in[0,T]$, respectively:
$(\widetilde{\mathcal{G}}_{net}^{i})_{t_{k}}=\begin{cases}1,&\text{if $t_{k}\in T_{peak}^{i}$},\\ 0,&\text{otherwise}.\end{cases}$ (17)
$(\widetilde{\mathcal{G}}_{bin}^{i})_{t_{k}}=\begin{cases}1,&\text{if $t_{k}\in T_{shift}^{i}$},\\ 0,&\text{otherwise}.\end{cases}$ (18)
@@
The code for implement our model and reproduce our analyses can be found at https://github.com/mauruscz/Gentrification.
Acknowledgements
@@
Supplementary Information
Supplementary Information 1: Binary Vector Transformation for Cross-Correlation Analysis
In our analysis, we compute the cross-correlation $R^{i}$ between two time series for a cell $i$. To facilitate this process, we transform both time series into binary “barcodes” that represent significant changes or peaks in the original signals. This section details the process of converting the original time series into their corresponding binary vectors, as illustrated in Figure S1.
The upper part of Figure S1 displays two example time series, denoted as $\mathcal{G}_{net}$ and $\mathcal{G}_{bin}$. We identify key events in these time series: peaks in the continuous signal $\mathcal{G}_{net}$ and transitions in the binary signal $\mathcal{G}_{bin}$. These events are then encoded as binary vectors, $\widetilde{\mathcal{G}}_{net}^{i}$ and $\widetilde{\mathcal{G}}_{bin}^{i}$, representing the time points at which the respective events occur. Specifically:
  • For $\widetilde{\mathcal{G}}_{net}^{i}$: A peak detection algorithm is applied to the time series $\mathcal{G}_{net}$. If a peak is detected at a specific time $t_{k}$, we assign $(\widetilde{\mathcal{G}}_{net}^{i})_{t_{k}}=1$; otherwise, $(\widetilde{\mathcal{G}}_{net}^{i})_{t_{k}}=0$. Thus, the binary vector $\widetilde{\mathcal{G}}_{net}^{i}$ captures the occurrence of peaks in $\mathcal{G}_{net}$ over the time period $T$.
  • For $\widetilde{\mathcal{G}}_{bin}^{i}$: Similarly, we track the 0-1 transitions in the binary signal $\mathcal{G}_{bin}$. If a transition from 0 to 1 occurs at time $t_{k}$, we assign $(\widetilde{\mathcal{G}}_{bin}^{i})_{t_{k}}=1$; otherwise, $(\widetilde{\mathcal{G}}_{bin}^{i})_{t_{k}}=0$. This forms the binary vector $\widetilde{\mathcal{G}}_{bin}^{i}$, encoding significant shifts in $\mathcal{G}_{bin}$.
The lower part of Figure S1 illustrates the resulting binary vectors, $\widetilde{\mathcal{G}}_{net}^{i}$ and $\widetilde{\mathcal{G}}_{bin}^{i}$, which are derived from the original time series. These binary representations, also referred to as ”barcodes,” simplify the identification of coinciding events (peaks and transitions) between the two time series, making it easier to compute the cross-correlation $R^{i}$.
It is important to note that in our analysis, the ”left” time series (i.e., the series corresponding to negative lags) represents $\mathcal{G}_{net}$, while the ”right” time series (i.e., the series corresponding to positive lags) represents $\mathcal{G}_{bin}$.
Figure S1: Example of barcode
Supplementary Information 2: Comparison with null model
We devised a null model out of our main model. In this version, agents behave as pure random walkers:
  • At each time step, every agent decides whether to move with a uniformly random probability (e.g., 50% chance of moving).
  • If an agent decides to move, it relocates to a cell with space chosen with a uniformly random probability throughout the grid.
The null model completely removes both the evaluation criteria and targeted relocation, providing a baseline for comparison with purely random motion. We applied the same statistical analysis used in the main text to assess the statistical relevance of $\mathcal{G}{net}$ anticipating $\mathcal{G}{bin}$ for this baseline model. The results are presented in Figure S2. Two key observations emerge:
  1. 1.
    The null model does not produce statistically significant results, indicating an absence of observable anticipatory patterns.
  2. 2.
    The peak of the cross-correlation, although not significant, occurs at $\tau=-1$. This is an expected outcome, as the $\mathcal{G}{net}$ measure is based on flow data, while $\mathcal{G}{bin}$ is derived from count data. Specifically, the flow of agents moving at one time step ($\mathcal{G}{net}$) is reflected in the count of agents ($\mathcal{G}{bin}$) in the subsequent time step, naturally creating a lag of -1.
The absence of statistically significant anticipatory patterns in the null model reinforces the validity of our main model’s findings. It suggests that the observed anticipatory behaviour is not a random artefact but rather an emergent property arising from the interplay of our model’s evaluation criteria and targeted relocation strategies. This comparison provides strong evidence for the robustness of our main model and the significance of its results in capturing complex socio-economic dynamics.
Figure S2: Cross-correlation analysis of $\mathcal{G}{net}$ and $\mathcal{G}{bin}$ for the null model. The x-axis represents the lag $\tau$, while the y-axis shows the cross-correlation coefficient. The shaded area indicates the 95% confidence interval. Note the lack of statistically significant anticipatory patterns in the null model, with peak at $\tau=-1$ due to the nature of the measured variables.
Supplementary Information 3: Early warning consistency across parameters and measures
In the main text, we presented the results of the cross-correlation analysis for a single value of $p_{H}$, the probability of high-income agents relocating. Here, we demonstrate that the observed anticipatory patterns are robust across different values of $p_{H}$ and alternative analytical measures.
Figure S3 illustrates the consistency of the cross-correlation analysis results for various $p_{H}$ values. The anticipatory relationship between $\mathcal{G}_{net}$ and $\mathcal{G}_{bin}$ remains stable across different probabilities of high-income agent relocation, indicating that this phenomenon is not sensitive to specific parameter choices within our model.
To further validate our findings, we conducted an additional analysis using Mutual Information (MI) instead of cross-correlation, as shown in Figure S4. MI provides a more general measure of statistical dependence, capturing both linear and non-linear relationships between variables. The results from this analysis corroborate our cross-correlation findings, exhibiting similar anticipatory patterns. It’s worth noting that while the overall trends are consistent, the absolute values of MI are approximately one order of magnitude smaller than those of the cross-correlation analysis.
The consistency of results across different $p_{H}$ values and analytical methods reinforces the robustness of our findings, suggesting that the anticipatory patterns observed in our model are a fundamental feature of the gentrification dynamics we’ve simulated, rather than an artifact of specific parameter choices or analytical approaches.
Figure S3: Cross-correlation analysis between $\mathcal{G}_{net}$ and $\mathcal{G}_{bin}$ for various $p_{H}$ values. Each panel represents a different $p_{H}$. The x-axis shows the lag $\tau$, and the y-axis represents the cross-correlation coefficient. Shaded areas indicate 95% confidence intervals. The consistent peak at negative lags across all $p_{H}$ values demonstrates the robustness of the anticipatory relationship between $\mathcal{G}_{net}$ and $\mathcal{G}_{bin}$, regardless of the probability of high-income agent relocation.
Figure S4: Mutual Information (MI) analysis between $\mathcal{G}_{net}$ and $\mathcal{G}_{bin}$ for various $p_{H}$ values. Each panel corresponds to a different $p_{H}$. The x-axis represents the lag $\tau$, while the y-axis shows the MI value. Shaded areas indicate 95% confidence intervals. The patterns observed are consistent with the cross-correlation analysis, exhibiting peaks at negative lags. Note that MI values are approximately one order of magnitude smaller than cross-correlation coefficients, which is expected due to the different nature of these measures. The consistency across $p_{H}$ values further supports the robustness of the anticipatory relationship between $\mathcal{G}_{net}$ and $\mathcal{G}_{bin}$.
Supplementary Information 4: Results are Consistent for a Larger (9x9) City
To assess the scalability and robustness of our findings, we extended our analysis to a larger urban environment, specifically a 9x9 grid city. This expansion allows us to verify whether the patterns observed in our original model persist in a more complex urban setting with a greater number of neighborhoods and potential interactions.
Figure S5 presents the spatial arrangement and income distribution for this larger city model.
Figure S5: Income distribution and spatial arrangement in a 9x9 city model. Left panel: Heatmap depicting the median income of each cell at the initial state of the simulation for a 9x9 grid city. Darker colors indicate higher median incomes. The hierarchical structure and spatial clustering of income levels are preserved in this larger model. Right panel: log-log plot of the initial distribution of agent incomes in the 9x9 city simulation, sampled from the 2022 USA Social Security Administration report. The distribution maintains the same shape as in the main model. N = $2^{12}$ agents.
Figure S6 presents the results of simulations conducted on the 9x9 city grid, mirroring the analysis shown in Figure 2 of the main text. Remarkably, the gentrification patterns observed in this larger model closely resemble those of the original smaller model, providing strong evidence for the scalability and consistency of our findings.
In the 9x9 city model, we observe that the critical role of $H$ agent mobility in initiating gentrification is preserved. At $p_{H}=0$, both $\mathcal{G}_{bin}^{city}$ and $\mathcal{G}_{net}^{city}$ show 0% gentrification, indicating a complete absence of the phenomenon when high-income agents are static. The abrupt transition to gentrification with minimal $H$ agent mobility ($p_{H}=0.01$) is also evident in the larger model, with both metrics showing gentrification in a significant proportion of cells. As $p_{H}$ increases, we see a similar monotonic rise in gentrification levels, with $\mathcal{G}_{bin}^{city}$ and $\mathcal{G}_{net}^{city}$ exhibiting comparable trends but slightly different magnitudes, consistent with the original model.
Figure S6: Gentrification dynamics in a 9x9 city model. We show gentrification levels and times across 150 simulations with fixed $\epsilon=20$, $N=2^{12}$, and varying values of $p_{H}$ (probability of high-income agent relocation). The left panel shows gentrification levels according to $\mathcal{G}_{bin}^{city}$, while the right panel displays levels according to the network-based measure $\mathcal{G}_{net}^{city}$. The x-axis represents different $p_{H}$ values, and the y-axis shows the percentage of gentrified cells. Box plots display the distribution of gentrification levels across simulations. The patterns observed closely resemble those in the smaller model: absence of gentrification at $p_{H}=0$, abrupt transition to gentrification at low $p_{H}$ values, and monotonic increase in gentrification levels as $p_{H}$ increases.
Figure S7 presents the results of simulations conducted on the 9x9 city grid, mirroring the analysis shown in Figure 5 of the main text. The gentrification patterns observed in this larger model closely resemble those of the original smaller model, providing strong evidence for the scalability and consistency of our findings.
In the bigger city model, we observe that the relationship between urban density and gentrification levels remains consistent across different values of the model parameter $p_{H}$. As in the smaller model, city-wise gentrification levels, captured by both $\mathcal{G}_{net}^{city}$ and $\mathcal{G}_{bin}^{city}$, show a clear trend: as city population density increases, so does the propensity for gentrification. This effect is amplified by the $H$ agents’ relocation rate $p_{H}$, with the curves increasing monotonically with $N$ for most values of $p_{H}$.
The relationship between city density and the average convergence time $\langle T\rangle$ in the larger model also mirrors the findings from the original model. The average convergence time increases with both $N$ and $p_{H}$, except when $p_{H}$ = 0. In extremely dense scenarios with stationary $H$ agents, the model may not reach the termination condition within the imposed step limit, consistent with the behavior observed in the smaller model.
Figure S7: Urban density and gentrification dynamics in a 9x9 city model. We show average results across 150 simulations for varying agent populations (x-axis, logarithmic scale) and high-income agent relocation rates $p_{H}$ (colors). Left panel: City-wise gentrification levels measured by $\mathcal{G}_{net}^{city}$. Middle panel: City-wise gentrification levels measured by $\mathcal{G}_{bin}^{city}$. Right panel: Average number of simulation time steps (convergence time $\langle T\rangle$). The patterns observed closely resemble those in the smaller 7x7 model: (1) gentrification levels increase with urban density for both measures, (2) this effect is amplified by higher $p_{H}$ values, and (3) convergence time generally increases with both density and $p_{H}$.
These observations suggest that the fundamental dynamics of our model, including the processes driving gentrification are not artifacts of the specific grid size used in the main text. Instead, they appear to be scalable properties that emerge from the underlying mechanisms of our agent-based model.

### v3

@@ 5 Discussion
We find that even a small proportion of high-income agents (5% of the population) significantly affect the dynamics of gentrification. When high-income agents do not move, even if still present in the model city, no gentrification is detected, and the model does not converge within the imposed limit of 300 steps. However, introducing even a low probability for the movement of high-income agents ($p_{H}$ = 1%) leads to model convergence in approximately 150 steps, while 40% of the city neighbourhoods experience at least one gentrification wave. Interestingly, when the model is initialized with a random configuration in which agents of all three types are substantially represented in each neighborhood (see Supplementary Figures 10–13), rather than with a core-periphery structure, it rapidly converges to a segregated state. This outcome is equivalent to the final configuration obtained in simulations starting from the core-periphery setup, although it lacks any preserved spatial organization.
@@
Our measures, $\mathcal{G}_{count}$ and $\mathcal{G}_{net}$, represent two distinct approaches to quantify gentrification. The count-based measure $\mathcal{G}_{count}$ evaluates the concentration of agents of the three types in each neighborhood over a time window $\Delta$. A significance threshold ($n_{H,M}^{*}$) is needed to detect when middle- and high-income agents are over-represented and, therefore, define the binarised version of the count-based measure, $\mathcal{G}_{bin}$. Such measure thus captures neighborhood transitions from under- to over-representation of middle- and high-income agents, indicating the completion of gentrification. The network-based measure $\mathcal{G}_{net}$ tracks the net inflow of middle- and high-income agents and the simultaneous outflow of lower-income agents over $\Delta$. This measure is thus rooted in temporal network analysis, where the existence of structures[24, 9, 42] in the networks under study is related to the simultaneity of the interactions (edges) between pairs or groups of nodes. Furthermore, this network-based approach avoids more or less arbitrary thresholds and aligns with available commuting flow data[4, 37, 40, 33], which could serve as a proxy given the absence of residential relocation records. $\mathcal{G}_{net}$ helps identifying early signs of gentrification by tracking peaks or plateaus in its trajectory, correlating higher-income resident influx with lower-income displacement. The cross-correlation computed between the count-based and the network-based measures highlights how peaks in $\mathcal{G}_{net}$ are consistently and significantly observed in advance with respect to $\mathcal{G}_{count}$. The analysis of the randomized versions of our model, in Supplementary Information 1, shows that while gentrification events caused by random agent-relocations are still detected, no significant cross-correlation exists between our two measures, emphasizing the importance of the agents’ decision-making rules in our model for predicting transitions. Our results hold true also when relaxing the termination condition of our model, thus allowing longer simulations (Supplementary Information 8). Furthermore, the analyses of longer simulations reveal a quasi-periodic occurrence of gentrification waves, defined as peaks of the time series of the volume of gentrification events unfolding across the whole city. These results demonstrate that our minimal model captures the essential features needed to reproduce gentrification: a heavy-tailed income distribution, few income classes, and distinct relocation strategies – notably the profit-driven behavior of high-income agents. Moreover, our network-based measure $\mathcal{G}_{net}$ enables earlier detection of gentrification compared to count-based metrics, potentially aiding policymakers in preventing low-income displacement.
@@
While our model provides valuable insights into gentrification dynamics, it has limitations that future research could address. The constant population size and static income-group assignments could be expanded to incorporate population growth and inter-city migration flows, potentially using a network-based approach to disentangle endogenous and exogenous causes of gentrification[21]. With the implementation of these extensions of our model, a complementary measure of impoverishment of a neighbourhood, such as the one we propose in the Supplementary Material (Supplementary Material 7) could help elucidate, in a quantitative manner, the causal relationships between the periods of impoverishment/disinvestment and gentrification of neighbourhoods. Multiple property ownership per agent could be introduced to model wealth concentration and short-term rental effects. The grid-based urban representation could be enhanced with more complex geographical features, although our results hold for both $\times7$ and $\times9$ grids (Supplementary Information 3). Moreover, simulating policy interventions[34] could provide insights for urban planners, particularly regarding density restrictions given our findings on city density and gentrification probability.
In conclusion, our agent-based model and novel quantitative measures offer a powerful framework for understanding and predicting gentrification processes. This approach not only advances our theoretical understanding of gentrification but also provides quantitative what-if tools for early detection and potential mitigation of its effects in real-world urban environments.
6 Methods
6.1 Agents’ income.
@@ 6 Methods
6.2 Gentrification: peaks, shifts and aggregate measure
@@
$T_{shift}^{i}=\{t:\mathcal{G}_{bin}^{i}(t,\Delta)=1\wedge\mathcal{G}_{bin}^{i}(t-1,\Delta)=0\}.$ (14)
@@
$T_{peak}^{i}=\left\{t:\left[\begin{array}[]{c}(\mathcal{G}_{net}^{i}(t,\Delta)>\mathcal{G}_{net}^{i}(t-1,\Delta)\wedge\mathcal{G}_{net}^{i}(t,\Delta)>\mathcal{G}_{net}^{i}(t+1,\Delta))\\ \vee\\ (\mathcal{G}_{net}^{i}(t,\Delta)>\mathcal{G}_{net}^{i}(t-1,\Delta)\wedge\mathcal{G}_{net}^{i}(t,\Delta)=\mathcal{G}_{net}^{i}(t+1,\Delta))\end{array}\right]\right\}.$ (15)
@@
$\mathcal{G}_{bin}^{city}=\frac{\sum_{i=1}^{N_{\text{cells}}}\mathbf{1}(|T_{shift}^{i}|>0)}{N_{\text{cells}}}\times 100\quad,\quad\mathcal{G}_{net}^{city}=\frac{\sum_{i=1}^{N_{\text{cells}}}\mathbf{1}(|T_{peak}^{i}|>0)}{N_{\text{cells}}}\times 100,$ (16)
@@
6.3 Lagged cross-correlation
We compute the mean cross-correlation $\langle R\rangle$ by calculating the lagged cross-correlation between pairs of time series of $\mathcal{G}^{i}_{net}(t,\Delta)$ and $\mathcal{G}^{i}_{bin}(t,\Delta)$ corresponding to each cell on the grid-view of a model city, for several values of the lag $\tau$, and then averaging over all cells. To compute the cross-correlation $R^{i}$ between the two time series for a cell $i$, we transform the two time series into two binary vectors (see Suplementary Information 4 for further details) $\widetilde{\mathcal{G}}_{net}^{i}$ and $\widetilde{\mathcal{G}}_{bin}^{i}$ of length $T$, where the $t_{k}$-th entry is $1$ if the original corresponding time series has a peak or 0-1 transition at time $t_{k}\in[0,T]$, respectively:
[…]
@@ 6.3 Lagged cross-correlation
The code for the implementation of our model and to reproduce our analyses can be found at https://github.com/mauruscz/Gentrification.
Acknowledgments
@@
ORCID